Betty’s Bakery has the following standard cost sheet for one unit of its most popular cake:

SQ SP
Direct materials 1.2 pounds $ 1.50 per pound
Direct labor 0.8 hours $ 12.00 per hour

During the month of May, the company made 920 cakes and incurred the following actual costs.

Direct materials purchased and used (1,380 pounds) $1,794

Direct labor (980 hours) $11,270

Required:

1. Calculate the direct materials price variance.

2. Calculate the direct materials quantity variance.

3. Calculate the direct materials variance.

4. Calculate the direct labor rate variance.

5. Calculate the direct labor efficiency variance.

6. Calculate the direct labor spending variance.

(For all requirements, Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)
Answer #1

1) Material price variance = (standard price - actual price) x actual quantity

= (1.50 - 1.3) x 1,380

= 276 F

2) Material quantity variance = (standard quantity - actual quantity) x standard price

= (1,104 - 1,380) x 1.5

= -414 U

3) direct material variance =  (Standard Quantity x Standard Price) -(Actual Quantity x Actual Price)

= (1,104 x 1.5) - (1,380 x 1.3)

= 1,656 - 1,794

= -138 U

4) Labour rate variance = (standard rate - actual rate) x actual hours used

= (12 - 11.5) x 980

= 490 F

5) Labour efficiency variance = (standard hour - actual hour) x standard rate

= (736 - 980) x 12

= -2,928 U

6) direct labour spending variance =  (Standard Rate x Standard Hour) – (Actual Rate x Actual Hour)

= (12 x 736) - (11.5 x 980)

= 8,832 - 11,270

= -2,438 U